OpenStudy (anonymous):

9. Rewrite this equation in standard form (Ax + By = C): y = -3x +15

OpenStudy (anonymous):

it can be y+3x=15

OpenStudy (anonymous):

or any multiple of this

OpenStudy (anonymous):

like 2y+6x=30etc

OpenStudy (anonymous):

just shift -3x frm Rhs to Lhs i mean

OpenStudy (anonymous):

no no nothing like that

OpenStudy (anonymous):

it is b here is 1 thats it

OpenStudy (anonymous):

no need by the way all u can do more is to div both side by 15 so u ll get (y/15)+(x/3)=1